# Hilbert-Kunz multiplicity of fiber products and idealization rings

### v3

@@ 5. Some bounds for the Hilbert-Kunz multiplicity
$R_{p,d}:=\frac{k[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
@@
    If $R$ is not regular, then $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})\geq 1+m_{d}$.
@@
    If $e_{\text{HK}}(R)=e_{\text{HK}}(R_{p,d})$, then the $\mathfrak{m}$-adic completion of $R$ is isomorphic to $R_{p,d}$ as local rings.
@@
Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$. The first inequality of part (1) of the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]. Also, recently, Trivedi [22] proved the second inequality of part (1) of the conjecture, for $d\geq 4$ and $p>d-3$.
In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular. The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas presented in the sections 3 and 4.
5.4 Proposition.
Let $d\geq 2$ be an integer and let $p>2$ be a prime number. Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p$ as in (2.5) and assume $R$ is a $d$-dimensional complete intersection, not regular. If $\dim R\geq\dim S>\dim T$, then
$e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the fiber product $R\times_{T}S$.
@@
The proof follows by Proposition 3.2, [11, Theorem 4.6] and [22, Theorem (A)]. ∎
@@
Let $d\geq 2$ be an integer, $p>2$ be a prime number, $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular, and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the amalgameted duplication $R\bowtie I$.
5.6 Proposition.
Let $d\geq 2$ be an integer, $p>2$ be a prime number and let $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular. If $M$ is a finite $R$-module, then
$e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the idealization ring $R\ltimes M$.
@@
The proof follows by item (1) of Corollary 4.3, [11, Theorem 4.6] and [22, Theorem (A)]
5.7 Remark.
Propositions 5.2 and 5.4, 5.6 effectively expand the class of rings and ideals that satisfy the desired inequalities of conjectures 5.1 and 5.3, respectively, because an idealization ring $R\ltimes M$ and a fiber product ring $R\times_{T}S$ may not even be Cohen-Macaulay when $R$ is a complete intersection local ring. Both situations are exemplified below.
5.8 Example ([3], Example 4.15).
Let $(R,\mathfrak{m})$ be an $d$-dimensional regular local ring with $x_{1},\dots,x_{d}$ a minimal basis for $\mathfrak{m}$. Let $M_{i}=R/(x_{1},\dots,x_{i})$, $0\leq i\leq d$ (for $i=0$, $M_{i}=R$). Then $\dim R\ltimes M_{i}=d$ and $\operatorname{depth}R\ltimes M_{i}=d-i$. So $R\ltimes M_{i}$ is not Cohen-Macaulay, for $0<i\leq d$.
5.9 Example.
Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].

### v4

@@ 6.2. Application: Watanabe-Yoshida Conjecture for Fiber Products and Idealization Rings
$A_{p,d}:=\frac{\overline{\mathbb{F}_{p}}[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
Let $(A,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
@@
    If $A$ is not regular, then $e_{\text{HK}}(A)\geq e_{\text{HK}}(A_{p,d})\geq 1+m_{d}$.
@@
    If $e_{\text{HK}}(A)=e_{\text{HK}}(A_{p,d})$, then the $\mathfrak{m}$-adic completion of $A$ is isomorphic to $A_{p,d}$ as local rings.
@@
This conjecture has a positive answer for local rings with dimension $\leq 7$ (see [10]), but for larger dimensions, the problem remains open. Some other positive answers in certain cases are provided, for instance, in [28, 13].
For the remainder of this section, we say that a $d$-dimensional Noetherian local ring $A$ of characteristic $p>0$ satisfies (WY1) provided that $A$ satisfies the strong inequality $e_{HK}(A)\geq e_{HK}(A_{p,d})$ of Watanabe-Yoshida conjecture 6.11(1). In our work, we obtained explicit formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and proved that these rings are not regular. We use these formulas to obtain new classes of singular rings that satisfy (WY1).
6.12 Theorem.
Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$. Assume that $\dim R=\dim R\times_{T}S$. If $R$ satisfies (WY1), then the fiber product $R\times_{T}S$ satisfies (WY1).
@@
By the explicit formulas given in Theorem 3.2, it is easy to see that this assertion holds when $\dim R\geq\dim S>\dim T$. Thus, we may assume that $\dim R=\dim S=\dim T$, and then $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. It is easy to see that $\pi_{S}(\mathfrak{m}_{S})=\mathfrak{m}_{T}$ and considering the exact sequence of $S$-modules
$0\rightarrow\operatorname{Ker}\pi_{S}\rightarrow S\xrightarrow{\pi_{S}}T\rightarrow 0,$
we have:
$e_{HK}(S)-e_{HK}(\mathfrak{m}_{S},T)$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})$
$e_{HK}(S)-\lim_{e\to\infty}\frac{\ell_{S}(T/\mathfrak{m}_{S}^{[q]}T)}{q^{\dim S}}$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})$
$e_{HK}(S)-e_{HK}(T)=e_{HK}(S)-\lim_{e\to\infty}\frac{\ell_{T}(T/\mathfrak{m}_{T}^{[q]}T)}{q^{\dim T}}$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})\geq 0.$
Thus, since $R$ satisfies (WY1), we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)\geq e_{HK}(A_{p,d})$. ∎
@@
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be a proper ideal of $R$. If $R$ satisfies (WY1), then the amalgamated duplication $R\bowtie I$ satisfies (WY1).
6.14 Theorem.
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. If $R$ satisfies (WY1), then the idealization $R\ltimes M$ satisfies (WY1).
@@
The assertion follows from the formula given in Corollary 4.3(1). ∎
An interesting consequence of Theorems 6.12 and 6.14 and Corollary 6.13 is due to a result of Enescu and Shimomoto [13, Theorem 4.6], which shows that every complete intersection ring (not regular) of dimension $d\geq 2$ and characteristic $p>2$ satisfies (WY1). It is important to observe that if $R$ is a complete intersection, the rings $R\times_{T}S$ and $R\ltimes M$ are not always complete intersections (see Examples 3.9 and 4.6). Therefore, the next result provides new classes of non-regular local rings that satisfy the strong inequality of Conjecture 6.11.
6.15 Corollary.
Let $d\geq 2$ and $p>2$. If $R$ is a $d$-dimensional complete intersection ring of characteristic $p$, not regular, then the following assertions hold:
  1. (1)
    If $d=\dim R\times_{T}S$, where $S,T$ are Noetherian local rings of characteristic $p$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(A_{p,d})$.
  2. (2)
    If $I$ is a proper ideal of $R$, then $e_{HK}(R\bowtie I)\geq e_{HK}(A_{p,d})$.
  3. (3)
    If $M$ is a finitely generated $R$-module, then $e_{HK}(R\ltimes M)\geq e_{HK}(A_{p,d})$.
Another interesting consequence of Theorems 6.12 and 6.14 and Corollary 6.13 is due to a result of Aberbach of Enescu [2, Section 4], which shows that the Watanabe-Yoshida Conjecture 6.11(1) holds if the ring $R$ has Hilbert-Samuel multiplicity $e(R)\leq 5$. It is important to observe that even if $R$ satisfies $e(R)\leq 5$, the rings $R\times_{T}S$ and $R\ltimes M$ may have Hilbert-Samuel multiplicity greater than five in several cases (e.g., if $\dim R=\dim S>\dim T$ and $\dim R=\dim M$, respectively). Therefore, the next result provides new classes of non-regular local rings that satisfy the strong inequality of Conjecture 6.11.
6.16 Corollary.
Let $d\geq 2$ and $p>2$. If $R$ is a $d$-dimensional Noetherian local ring of characteristic $p$, not regular, with $e(R)\leq 5$, then the following assertions hold:
  1. (1)
    If $d=\dim R\times_{T}S$, where $S,T$ are Noetherian local rings of characteristic $p$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(A_{p,d})$.
  2. (2)
    If $I$ is a proper ideal of $R$, then $e_{HK}(R\bowtie I)\geq e_{HK}(A_{p,d})$.
  3. (3)
    If $M$ is a finitely generated $R$-module, then $e_{HK}(R\ltimes M)\geq e_{HK}(A_{p,d})$.
Acknowlegment.
The first author was supported by Sao Paulo Research Foundation (FAPESP) under grant 2019/21181-0. The second author was supported by Sao Paulo Research Foundation (FAPESP) under grant 2022/12114-0.